### v2

# Slab tilings, flips and the triple twist

Abstract
A domino is a $2\times 1\times 1$ parallelepiped formed by the union of two unit cubes and a slab is a $2\times 2\times 1$ parallelepiped formed by the union of four unit cubes. We are interested in tiling regions formed by the finite union of unit cubes. Domino tilings have been studied before; here we investigate slab tilings. As for domino tilings, a flip in a slab tiling is a local move: two neighboring parallel slabs are removed and placed back in a different position.
Inspired by the twist for domino tilings, we construct a flip invariant for slab tilings: the triple twist, assuming values in ${\mathbb{Z}}^{3}$. We show that if the region is a large box then the triple twist assumes a large number of possible values, roughly proportional to the fourth power of the volume. We also give examples of smaller regions for which the set of tilings is connected under flips, so that the triple twist assumes only one value.
00footnotetext: 2010 Mathematics Subject Classification. Primary 05B45; Secondary 52C20, 52C22, 05C70. Keywords and phrases Three-dimensional tilings, dominoes, dimers, slabs
1 Introduction
In this paper we consider tilings of cylinders, of which rectangular boxes are a special case. A cylinder ${\cal R}$ is a cubiculated region in $\mathbb{R}^{3}$ of the form ${\cal R}={\cal D}\times[0,N]$, where ${\cal D}$ is a disk, a quadriculated connected and simply connected region on the $z=0$ plane. Unit cubes are of the form $[x,x+1]\times[y,y+1]\times[z,z+1]$ where $(x,y,z)\in{\mathbb{Z}}^{3}$.
We give a brief summary of the directly relevant results about 2d and 3d domino tilings. We choose to depict a three-dimensional domino tiling by floors, numbered from left to right. The $n$-th floor from left to right corresponds to $z\in[n,n+1]$, $n\in{\mathbb{Z}}$. A domino which is contained in only one floor (named horizontal domino) is represented as a $2\times 1$ light grey rectangle. A domino which crosses two floors (named vertical) is represented as two $1\times 1$ squares which have the same projection on the first floor, with the square on the lower floor (left) being painted dark grey and the square on the upper floor (right) being painted white. Figure 1 shows a domino tiling of the $3\times 3\times 2$ box. On the left, the floor $z\in[0,1]$; on the right, the floor $z\in[1,2]$. The $x$ and $y$ axes are usually drawn as in this example. For convenience, we may at times speak of a region using vocabulary of the floor diagram for the region. That is, we talk of the corresponding square of the floor diagram when referring to a certain cube.
$y$
$x$
Figure 1: Domino tiling of a $3\times 3\times 2$ box
We are interested in local moves. The most natural local move is a flip: remove two parallel dominoes and place them back in the only other possible position. We provide a graph theoretical formulation of this scenario. Given a region ${\cal D}$, construct a new graph, where vertices are domino tilings ${\mathbf{t}}$ of ${\cal D}$ and edges join pairs ${\mathbf{t}}_{0},{\mathbf{t}}_{1}$ of tilings which differ by a flip. If ${\cal D}\subset{\mathbb{R}}^{2}$ is a quadriculated disk, i.e., a bounded 2d contractible region, then this graph is connected. Equivalently, any two tilings of ${\cal D}$ can be joined by a finite sequence of flips [8]. Applied to other regions (possibly 3d), this construction allows us to speak of connected components under flips.
For 3d domino tilings, this new graph is almost never connected. As a first example, consider the box in Figure 1 and the graph whose vertices are its tilings. This box admits $229$ tilings (see [6], which includes larger examples). The tiling in Figure 1 admits no flips (it corresponds to an isolated vertex in the graph). The mirror image is likewise isolated. The other $227$ tilings form a single connected component under flips. It might be conjectured after this example that almost all tilings of a 3d box can be joined by a finite number of flips: as we shall see, this is also not correct.
Given a contractible cubiculated region ${\cal R}\subset{\mathbb{R}}^{3}$, there exists a map $\operatorname{Tw}$ (twist) from the set of domino tilings of ${\cal R}$ to ${\mathbb{Z}}$. There are several ways to construct the function $\operatorname{Tw}$: see [1, 3] for more general constructions. For completeness, we include the construction from [4] in Section 2: that construction is valid if ${\cal R}$ is a cylinder. This map has several important properties. The main property is that twist is an invariant under domino flips, so that connected components under flips are contained in level sets of $\operatorname{Tw}$. For the $3\times 3\times 2$ box, the tiling in Figure 1 has $\operatorname{Tw}=-1$; its mirror image has $\operatorname{Tw}=+1$. The other $227$ tilings all have $\operatorname{Tw}=0$.
Twist also works well with symmetries of the region ${\cal R}$. For instance, with our construction, if ${\cal R}$ is a cylinder then the twist is left unchanged after a rotation in ${\mathbb{R}}^{3}$. Also for cylinders, the twist changes sign under reflections by any of the planes $x=0$, $y=0$ or $z=0$.
For domino tilings, there is another important local move besides the flip: the trit ([4, 1]). In order to perform a trit in a tiling ${\mathbf{t}}$, search for three dominoes whose union is a $2\times 2\times 2$ cube minus two unit cubes in opposite vertices. The trit is then performed by removing the three dominoes and placing them back in the only other possible position. A trit always changes the twist by adding ±$\pm 1$, where the sign depends on orientation. The tiling in Figure 1 admits four possible trits, all of them taking us to tilings of twist 00.
There is no fixed finite set of local moves connecting all domino tilings which works for any contractible cubiculated region ${\cal R}\subset{\mathbb{R}}^{3}$. It is conjectured that any two tilings of a box can be connected by a finite sequence of flips and trits. It is known for large classes of boxes that this holds for almost all tilings [6]. It is also known that any tiling of a box admits at least one flip or trit [2].
It is conjectured that for large boxes the twist follows a normal distribution. This is known to be true for boxes of size $L\times M\times N$ where $L,M\geq 4$ are even and fixed and $N$ tends to infinity [7]. In this case, it is also known that for almost all pairs $({\mathbf{t}}_{0},{\mathbf{t}}_{1})$ of tilings with $\operatorname{Tw}({\mathbf{t}}_{0})=\operatorname{Tw}({\mathbf{t}}_{1})$ there exists a finite sequence of flips connecting ${\mathbf{t}}_{0}$ and ${\mathbf{t}}_{1}$. This implies that connected components under flips almost coincide with level sets of the function $\operatorname{Tw}$. Equivalently, within a (large) level set of $\operatorname{Tw}$ there exists a giant component.
The aim of the present paper is to investigate related questions for slab tilings of 3d regions. A slab is a $2\times 2\times 1$ cuboid. Let us draw a small example in Figure 2. As for domino tilings, we depict a slab tiling by floors, numbered from left to right. A slab which is contained in only one floor (horizontal) is represented as a $2\times 2$ light grey square. A slab which crosses two floors (vertical) is represented as two planar dominoes. These two dominoes must have the same projection on the first floor. The domino on the lower (left) floor is painted dark grey and the domino on the upper (right) floor is painted white.
Figure 2: A slab tiling of a $6\times 6\times 6$ box
We define flips for slabs in the same way as for dominoes: we take two slabs which have a $2\times 2$ face in common, remove them from the region, and then place them back in one of the two other positions.
A natural question is to investigate connected components via flips of the set of slab tilings of a given region. The existence of the twist as an invariant under flips is crucial for the study of similar questions for domino tilings: does a similar invariant exist for slab tilings? As we shall see, the answer is yes, but there are many significant differences. The new invariant is the triple twist ($\operatorname{TTw}$), which assumes values in ${\mathbb{Z}}^{3}$. Its construction, which takes a large portion of the present paper, shows that it is related to the twist (for dominoes) and indeed invariant under flips. The triple twist assumes a large number of values.
Theorem 1.
Consider a cubiculated box ${\cal R}$ of size $N\times N\times N$, $N\in 2{\mathbb{N}}$. If $N$ is sufficiently large then there exists positive constants $C_{-}<C_{+}$ with the following properties. If ${\mathbf{t}}$ is a slab tiling of ${\cal R}$ then $|\operatorname{TTw}({\mathbf{t}})|<C_{+}N^{4}$. If $t\in(2{\mathbb{Z}})^{3}$ and $|t|<C_{-}N^{4}$ then there exists a tiling ${\mathbf{t}}$ of ${\cal R}$ with $\operatorname{TTw}({\mathbf{t}})=t$.
The previous theorem gives a lower bound for the number of connected components under flips of the set of slab tilings of a large box (see Proposition 23 for an explicit formula). We easily obtain an upper bound for the set of values of $\operatorname{TTw}$, but the number of connected components appears to be much larger (and harder to estimate). In some simple cases, such as a $4\times 4\times N$ or an $M\times N\times 2$ box, there is only one connected component.
The construction of the triple twist leads us to consider mixed tilings, consisting of horizontal slabs and vertical dominoes. Given a slab tiling, we construct a corresponding mixed tiling (Section 3). If the initial region is a box, this construction can be performed in any of the three canonical directions. Given in turn a mixed tiling, we use a four color pattern to obtain a corresponding domino tiling (Section 4): we then compute its twist. The triple twist is computed by performing the above construction in all three directions. As for the domino tilings, these new versions of the twist are well behaved under symmetries.
As for domino tilings, there exist examples of slab tilings which admit no flip (see Figures 20 and 21). There are similar examples of mixed tilings which admit no flip (see Figure 7). We find it more surprising to observe that there exists no fixed set of local moves which makes the set of slab tilings connected under such moves. More than that: given a fixed set of local moves (involving a bounded number of slabs) there exist slab tilings which admit none of these local moves (see Lemma 6.3, Corollary 6.4 and Remark 6.5).
In Section 2 we define and review basic properties of the twist for domino tilings. We introduce mixed tilings in Section 3. In Section 4 we consider a way of using four colors to paint the unit cubes in order to construct the twist invariants for mixed and slab tilings. In Section 5 we discuss some properties of these invariants, including the effect of symmetries. Section 6 includes examples and results concerning connectivity under local moves. Theorem 1 and related results are proved in Section 7. Finally, in Section 8 we list some open problems.
2 Construction of the twist
This section contains one among several known constructions of the twist for domino tilings of a cubiculated region ${\cal R}\subset{\mathbb{R}}^{3}$. This is included for completeness; we follow [4]. See [1, 3] for alternative constructions, sometimes more general.
The effect is a function which takes two dominoes in a domino tiling ${\mathbf{t}}$ and returns a value in $\{0,\pm 1\}$. On ${\mathbf{t}}$, we choose a horizontal domino $d_{0}$ such that the line through the centers of two cubes covered by $d_{0}$ is parallel to the base vector $\mathbf{e_{2}}$. Let $d_{1}$ be a vertical domino in ${\mathbf{t}}$. We say $d_{1}$ affects $d_{0}$ if any of the cubes of $d_{1}$ is on the same floor as $d_{0}$ and if the projections of the interiors of $d_{0}$ and $d_{1}$ over the plane $x=0$ have non-empty intersection. Figure 3 shows some examples for which there are effects. If these conditions do not hold we say the effect of $d_{1}$ on $d_{0}$ is 00.
In the case $d_{1}$ does affect $d_{0}$, we define the effect as a sign based on some parameters. The first is whether $d_{1}$ is on the left or on the right side of $d_{0}$. The second is whether $d_{1}$ is above or below $d_{0}$. The third is whether the distance between $d_{1}$ and $d_{0}$ is odd or even. And the fourth and final is whether $d_{1}$’s lower half cube is on or below $d_{0}$’s floor. As a convention, the effect on the first example of Figure 3 is set as $+1$, and a change on any of the parameters above yields a change of sign on the effect.
$y$
$x$
$+1$
$+1$
$-1$
$-1$
$-1$
Figure 3: Examples of effects for pairs of dominoes
We then define the twist $\operatorname{Tw}({\mathbf{t}})$ as $\frac{1}{4}$ times the sum of all the effects of $d_{1}$ on $d_{0}$ for all the suitable pairs $(d_{0},d_{1})$. If ${\cal R}$ is a cylinder and ${\mathbf{t}}$ is a tiling of ${\cal R}$ then $\operatorname{Tw}({\mathbf{t}})$ is an integer [4].
For example, the twist of the tiling in Figure 1 is $-1$, since each floor has a contribution of $-\frac{1}{2}$. If we appy a trit to this tiling, the twist of the new tiling is 00. In Figure 4 we show two examples of tilings of cylinders, and in Figure 5 we rotate each of them, noting the twist is unchanged, consistently with the remarks in the Introduction.
Figure 4: Two examples of tilings of cylinders; in both cases, $\operatorname{Tw}=0$
Figure 5: The $x$-view of the two tilings in Figure 4, i.e., the same tilings after a rotation; we still have $\operatorname{Tw}=0$
3 Mixed tilings
The study of slab tilings led us to consider another class of tilings: mixed tilings. Given a three dimensional region ${\cal R}\subset{\mathbb{R}}^{3}$ with a specified vertical direction, a mixed tiling is a decomposition of ${\cal R}$ into horizontal slabs and vertical dominoes. We admit as mixed tilings the configurations where there are only horizontal slabs or only vertical dominoes.
Figure 6: A non trivial example of mixed tiling
The mixed tiling in Figure 6 can be obtained from the slab tiling in Figure 2 by slicing each vertical slab in half along a plane perpendicular to the paper. This process of slicing vertical slabs in half to transform a slab tiling into a mixed tiling can always be done. However, it is not always possible to undo this process in a well defined manner. For example, in Figure 6 there are many ways to join the vertical dominoes which cross the first and second floor. Also, there are mixed tilings which do not come from any slab tiling, as in Figure 7.
Figure 7: A mixed tiling which does not correspond to a slab tiling, obtained from the tiling in Figure 6 by taking the central $4\times 4\times 6$ box
We can also define a flip in a mixed tiling. In one way, we take two adjacent slabs which lie on different floors, perform a slab flip and then slice the resulting vertical slabs in half. In the other way, we take four vertical dominoes which form a $2\times 2\times 2$ cube, join two pairs of adjacent dominoes and perform a slab flip to end with two horizontal slabs. The mixed tiling in Figure 7 admits no flip; See Figure 8 for an example of a flip.
Figure 8: We show one flip performed on the tiling of Figure 6
4 Coloring and twist
The goal of this section is to describe the process of transforming a slab or mixed tiling ${\mathbf{t}}$ of a given cylinder ${\cal R}$ into a domino tiling $\tilde{{\mathbf{t}}}$ of another cylinder $\tilde{{\cal R}}$. This will be used in the end to define the twist of a mixed tiling and the triple twist of a slab tiling.
The first step of this process is to consider a certain coloring of the cubes which compose our cylinder. The coloring uses four colors and must meet the condition that each slab covers exactly one cube of each color. Provisionally, this condition will be called slab-type. We use the colors green, blue, red and yellow.
$x$
$y$
Figure 9: A slab-type coloring of a $6\times 6\times 6$ box, based on Example 4.1
Example 4.1.
Here we give an algebraic description of a slab-type coloring. Let $\phi:{\mathbb{Z}}\oplus{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}/(2)\oplus{\mathbb{Z}}/(2)$ be the additive group homomorphism $\phi(x,y,z)=(y+z,x+z)$. Associate with each element of ${\mathbb{Z}}/(2)\oplus{\mathbb{Z}}/(2)$ a color and paint the cube $[x,x+1]\times[y,y+1]\times[z,z+1]$ with the color $\phi(x,y,z)$. As in Figure 9, $(0,0)$ is green, $(1,0)$ is yellow, $(0,1)$ is blue and $(1,1)$ is red. $\diamond$
Lemma 4.2.
Let ${\cal R}=[0,L]\times[0,M]\times[0,N]$ be a box with $\min\{L,M,N\}\geq 2$. There is, up to permutation of the colors, only one slab type coloring.
Proof.
Refer to Figure 10. Consider a slab which covers the yellow, red, $Z$ and $W$ squares. This means $W$ must be painted either blue or green. Now consider a slab which covers the blue, red, $Y$ and $W$ squares. The only possibility left is for $W$ to be painted green. Analogously, $X$ must be red, $Y$ must be yellow and $Z$ must be blue.
In a similar fashion, we rule out possibilities to propagate the painting to the remaining part of ${\cal R}$. ∎
$X$
$Y$
$Z$
$W$
Figure 10: The top left corner of the first two floors of ${\cal R}$
Remark 4.3.
The uniqueness property in Lemma 4.2 also holds for other kinds of regions and does not hold for some others. We do not think a more precise statement would be useful here.
Given Lemma 4.2, there is limited interest in considering more general colorings if, as in the present paper, our focus is on boxes. From now on, we will consider only the coloring in Example 4.1. $\diamond$
The second step of the process is to transform the region. We first consider cylinders of the form ${\cal R}={\cal D}\times[0,N]$ and the $z$-view. A pair of colors $(c_{0},c_{1})$ is good (for this view) if $c_{0}$ and $c_{1}$ are on a same diagonal on the first floor: the good pairs are therefore red-green and blue-yellow. The colors in a good pair are also called good; the other colors are bad. Informally, we keep all cubes (or squares) of one of the two good colors and delete all cubes of bad colors. We then inflate the remaining cubes so as to obtain cuboids which fill in the space left by the deleted cubes. Figure 11 shows an example: the initial region is as in Figure 9 and the good pair is red-green.
Figure 11: The region in Figure 9 is transformed by deleting the blue and yellow cubes and inflating the red and green ones
A more precise construction is in order. For the red-green pair, the cube $[x,x+1]\times[y,y+1]\times[z,z+1]$ is good if $x+y$ is even. In the notation of Example 4.1, the cube is good if $\phi(x,y,z)\in\{(0,0),(1,1)\}\subset{\mathbb{Z}}/(2)\oplus{\mathbb{Z}}/(2)$. If good, the above cube inflates to become the cuboid $\tilde{S}\times[z,z+1]$ where $\tilde{S}\subset{\mathbb{R}}^{2}$ is the square with vertices $(x-\frac{1}{2},y+\frac{1}{2})$, $(x+\frac{1}{2},y-\frac{1}{2})$, $(x+\frac{3}{2},y+\frac{1}{2})$, $(x+\frac{1}{2},y+\frac{3}{2})$. The region $\tilde{R}$ is the union of the resulting inflated cuboids, as in Figure 11. The new cuboids keep the old colors, so that $\tilde{R}$ is a cubiculated region. Strictly speaking, we might feel the need to rotate and rescale, but we usually refrain from doing so as it only complicates the description.
We are now left with the task of transforming the tilings. We first consider slab tilings; let ${\mathbf{t}}$ be a tiling of ${\cal R}$. Each slab $P$ of ${\mathbf{t}}$ covers two preserved cubes and two deleted cubes. Let $S_{0}$ and $S_{1}$ be the two preserved cubes, which are transformed into $\tilde{S}_{0}$ and $\tilde{S}_{1}$, respectively. We say that $P$ is taken to the domino $\tilde{P}$, the union of $\tilde{S}_{0}$ and $\tilde{S}_{1}$. See Figures 2 and 12 for an example. In the second row of Figure 12, we show the tiling rotated clockwise by $45^{\circ}$; in this example, we have $\operatorname{Tw}=+2$.
Figure 12: The domino tiling obtained from the slab tiling in Figure 2 or the mixed tiling in Figure 6
For mixed tilings the construction is similar. Horizontal slabs are transformed into horizontal dominoes, as above. Vertical dominoes consist of two cubes which are either both preserved or both removed. In the first case, the domino in transformed into a vertical domino; in the second case the domino is deleted; see Figures 6 and 12.
Now we consider ${\mathbf{t}}$ a slab tiling of a box. Until this moment, we have only looked at our tiling ${\mathbf{t}}$ of ${\cal R}$ from a particular position of the axes, as stated at the beginning of the section. This position of the axes is exactly that of an observer looking from the positive part of the $z$-axis. We say we are looking at the $z$-view ${\mathbf{t}}_{z}$ of ${\mathbf{t}}$. What we can actually do is look at the region ${\cal R}$ from the positions of the positive parts of the $x$ and $y$ axes, yielding the views ${\mathbf{t}}_{x}$ and ${\mathbf{t}}_{y}$, respectively. Strictly speaking, ${\mathbf{t}}_{x}$, ${\mathbf{t}}_{y}$ and ${\mathbf{t}}_{z}$ are three different views of the same tiling (as sets of slabs). What does change though is the floor diagram.
$z$
$x$
Figure 13: The figures corresponding to the $y$-view, with the blue-green good pair; $\operatorname{Tw}(\tilde{\mathbf{t}})=0$
$y$
$z$
Figure 14: The figures corresponding to the $x$-view, with the green-yellow good pair; $\operatorname{Tw}(\tilde{\mathbf{t}})=0$
For each view there are two good pairs of colors. For the $y$-view, the good pairs are green-blue and red-yellow; for the $x$-view, red-blue and green-yellow. In general, a pair of colors $\kappa$ defines an axis and an axis defines two complementary pairs of colors, $\kappa$ and $\bar{\kappa}$.
Given a view, a good pair of colors and a tiling of a box, we perform the above construction. In Figures 13 and 14 we continue the example of the slab tiling of the $6\times 6\times 6$ box given in Figure 2. Notice that the resulting regions and domino tilings $\tilde{\mathbf{t}}$ are quite different. When necessary, we write ~$\tilde{\cal R}_{\kappa}$ and $\tilde{\mathbf{t}}_{\kappa}$ to indicate the pair of colors.
We are ready to state the definitions of the twist.
Figure 15: Two possibilities of flipping two slabs on the top left corner of the first floor of Figure 2; under the transformation, the first possibility yields the same domino tiling as in Figure 12
Figure 16: The domino tiling from the second possibility in Figure 15 after the transformation
Definition 4.4 (Mixed twist).
Let ${\cal R}$ be a $z$-cylinder and $\kappa$ be a good pair of colors. Let ${\mathbf{t}}$ be a mixed tiling of ${\cal R}$. Consider the transformed region ~$\tilde{\cal R}_{\kappa}$ and the corresponding domino tiling $\tilde{{\mathbf{t}}}_{\kappa}$. The twist of ${\mathbf{t}}$ is
$\operatorname{Tw}_{\kappa}({\mathbf{t}})=\operatorname{Tw}(\tilde{{\mathbf{t}}}_{\kappa})\in\mathbb{Z}.$
Definition 4.5 (Triple twist).
Let ${\cal R}$ be a box and let ${\mathbf{t}}$ be a slab tiling of ${\cal R}$. Let $\kappa_{x},\kappa_{y},\kappa_{z}$ be good pairs of colors for the $x$, $y$ and $z$ views, respectively. Consider the transformed domino tilings $\tilde{\mathbf{t}}_{\kappa_{x}}$, $\tilde{\mathbf{t}}_{\kappa_{y}}$ and $\tilde{\mathbf{t}}_{\kappa_{z}}$. Let $\operatorname{Tw}_{\kappa}({\mathbf{t}})=\operatorname{Tw}(\tilde{\mathbf{t}}_{\kappa})$. The triple twist of ${\mathbf{t}}$ is
$\operatorname{TTw}_{\kappa_{x},\kappa_{y},\kappa_{z}}({\mathbf{t}})=(\operatorname{Tw}_{\kappa_{x}}({\mathbf{t}}),\operatorname{Tw}_{\kappa_{y}}({\mathbf{t}}),\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}}))\in\mathbb{Z}^{3}.$
5 Properties of the Triple Twist
The twist for domino tilings has several simple properties and some of them are inherited by the triple twist.
Let ${\mathbf{t}}$ be a slab or mixed tiling of a cylinder ${\cal R}$. Fix a good pair of colors $\kappa$. When performing a flip in ${\mathbf{t}}$, one of two things happens with the dominoes in the corresponding domino tiling $\tilde{\mathbf{t}}$: either a domino flip is performed or nothing changes at all. Since the domino twist is flip invariant, this means that the several $\operatorname{Tw}_{\kappa}$ defined above are invariants under flips.
The complement $\bar{\kappa}$ of a good pair of colors $\kappa$ is also good. One of the first questions which arises when defining the triple twist is regarding the choice of $\bar{\kappa}$ instead of $\kappa$. For example, the calculation of the twist in Figure 12 depended on the choice of the green-red pair in Figure 9. We could have chosen instead the pair blue-yellow. This other choice would even yield a different domino tileable region, but does not actually provide any new information, which is the content of Lemma 5.1.
Lemma 5.1.
Let ${\cal R}$ be a cylinder. Let $\kappa$ and $\bar{\kappa}$ be complementary good pairs of colors. Let ${\mathbf{t}}$ be a mixed tiling of ${\cal R}$. Then, $\operatorname{Tw}_{\kappa}({\mathbf{t}})=-\operatorname{Tw}_{\bar{\kappa}}({\mathbf{t}})$.
Proof.
Assume that the two good pairs are $\kappa$ for red-green and $\bar{\kappa}$ for yellow-blue. Throughout the proof, when talking about the overlap of two pieces, we refer to the overlap of their projections onto the $xy$ plane. Firstly, we take a look at how two horizontal slabs in different floors overlap. There are three possibilities: they can overlap in either one, two or four squares. As an illustrative example, we show these three cases in the mixed tiling of the $4\times 4\times 4$ box in Figure 17.
Figure 17: An example of a mixed tiling; the top left cube on the first floor is $[0,1]^{3}$, which is green
Figure 18: First tiling is red-green and second tiling is blue-yellow
We transform the mixed tiling ${\mathbf{t}}$ into two domino tilings $\tilde{\mathbf{t}}_{\kappa}$ and $\tilde{\mathbf{t}}_{\bar{\kappa}}$. In Figure 18 we show these two tilings for the example in Figure 17. We will calculate the two twists side by side, but in order to do so, we will rotate both regions by $45$ degrees as in Figure 4 and then look at them from the $x$-axis view, as depicted in Figure 5. By doing so, dominoes which were vertical in Figure 17 (and thus vertical in Figure 18) become horizontal in Figure 4 and hence will not affect the calculation of the twists.
In Figure 19 we look at the relative position of a pair of dominoes which comes from a pair of overlapping slabs (on different floors) in ${\mathbf{t}}$ after performing the sequence of transformations in the above paragraph. We do not require the pair of slabs to be on consecutive floors.
Figure 19: We have, in order: one instance of case 1, two instances of case 2 and two instances of case 3
  1. 1.
    If a pair of slabs overlaps in one square, the corresponding dominoes are parallel in both the green-red and blue-yellow tilings. In one of the tilings the dominoes will both be vertical with no overlap and in the other tiling both will be horizontal with an overlap of one square.
  2. 2.
    If a pair of slabs overlaps in four squares, the corresponding dominoes will be parallel in both the green-red and blue-yellow tilings, but in the two cases they will be horizontal and have an overlap of two squares.
  3. 3.
    If a pair of slabs overlaps in two squares, one of the corresponding dominoes will be vertical and the other horizontal, overlapping on one square. But in the green-red tiling the contribution to the twist will be $+1$ and on the blue-yellow tiling the contribution will be $-1$ (or vice-versa).
Therefore, summing the contributions of the twists, which are only nontrivial in case 3, we see that, indeed, the statement holds. ∎
We now consider the effect of spatial symmetries on the twist of mixed tilings. The reader should compare the next result with the corresponding simpler situation for twists of domino tilings.
Proposition 5.2.
Consider the box ${\cal R}=[-L,L]\times[-M,M]\times[-N,N]$, where $L,M,N$ are positive integers. Let ${\mathbf{t}}$ be a mixed tiling of ${\cal R}$, with vertical dominoes in the $z$ direction. Let $\kappa$ be a good pair of colors for the $z$-view.
  1. 1.
    Let $r_{x}$, $r_{y}$ and $r_{z}$ denote the reflections about the planes $x=0$, $y=0$ and $z=0$, respectively. We have
    $\operatorname{Tw}_{\kappa}(r_{x}({\mathbf{t}}))=\operatorname{Tw}_{\kappa}({\mathbf{t}}),\quad\operatorname{Tw}_{\kappa}(r_{y}({\mathbf{t}}))=\operatorname{Tw}_{\kappa}({\mathbf{t}}),\quad\operatorname{Tw}_{\kappa}(r_{z}({\mathbf{t}}))=-\operatorname{Tw}_{\kappa}({\mathbf{t}}).$
  2. 2.
    Assume that $L=M$ and let $\rho$ denote a $90^{\circ}$ rotation around the $z$ axis. We have $\operatorname{Tw}(\rho({\mathbf{t}}))=-\operatorname{Tw}({\mathbf{t}})$.
Proof.
We first consider case (item) 1 for $r_{x}$. Let $\kappa$ and $\bar{\kappa}$ be the red-green and blue-yellow pairs, respectively. Let ${\mathbf{t}}_{0}$ and ${\mathbf{t}}_{1}=r_{x}({\mathbf{t}}_{0})$ be mixed tilings. Let $\tilde{\mathbf{t}}_{i,\kappa}$ and $\tilde{\mathbf{t}}_{i,\bar{\kappa}}$ be the transformed tilings obtained from ${\mathbf{t}}_{i}$ under each good pair. The set of red-green cubes is taken by $r_{x}$ to the set of blue-yellow cubes. Thus, the reflection $r_{x}$ takes $\tilde{\mathbf{t}}_{i,\kappa}$ to $\tilde{\mathbf{t}}_{1-i,\bar{\kappa}}$ so that $\operatorname{Tw}(\tilde{\mathbf{t}}_{1-i,\bar{\kappa}})=-\operatorname{Tw}(\tilde{\mathbf{t}}_{i,\kappa})$. By Lemma 5.1 we have $\operatorname{Tw}(\tilde{\mathbf{t}}_{i,\bar{\kappa}})=-\operatorname{Tw}(\tilde{\mathbf{t}}_{i,\kappa})$ and therefore
$\operatorname{Tw}_{\kappa}({\mathbf{t}}_{1})=\operatorname{Tw}(\tilde{\mathbf{t}}_{1,\kappa})=-\operatorname{Tw}(\tilde{\mathbf{t}}_{1,\bar{\kappa}})=\operatorname{Tw}(\tilde{\mathbf{t}}_{0,\kappa})=\operatorname{Tw}_{\kappa}({\mathbf{t}}_{0}),$
as desired. Case 1 for $r_{y}$ is similar. The set of red-green cubes is taken to itself by $r_{z}$ and therefore, ${\mathbf{t}}_{3}=r_{z}({\mathbf{t}}_{0})$, case 1 for $r_{z}$ boils down to
$\operatorname{Tw}_{\kappa}({\mathbf{t}}_{3})=\operatorname{Tw}(\tilde{\mathbf{t}}_{3,\kappa})=-\operatorname{Tw}(\tilde{\mathbf{t}}_{0,\kappa})=-\operatorname{Tw}_{\kappa}({\mathbf{t}}_{0}),$
completing item 1. Finally, the rotation $\rho$ takes the set of red-green cubes to the set of blue-yellow cubes. Set ${\mathbf{t}}_{4}=\rho({\mathbf{t}}_{0})$: $\tilde{\mathbf{t}}_{4,\bar{\kappa}}$ is a rotation of $\tilde{\mathbf{t}}_{0,\kappa}$ so that $\operatorname{Tw}(\tilde{\mathbf{t}}_{4,\kappa})=-\operatorname{Tw}(\tilde{\mathbf{t}}_{4,\bar{\kappa}})=-\operatorname{Tw}(\tilde{\mathbf{t}}_{0,\kappa})$, completing the proof. ∎
The following proposition is similar, but for slab tilings. A symmetry of ${\mathbb{R}}^{3}$ which preserves unit cubes is a map of the form $q(w)=\mathbf{v_{0}}+\mathbf{Qw}$ where $\mathbf{v_{0}}\in{\mathbb{Z}}^{3}$ and $\mathbf{Q}$ is a $3\times 3$ permutation matrix with signs. Let $\operatorname{sign}(q)=\det(\mathbf{Q})\in\{\pm 1\}$.
Proposition 5.3.
Consider $q$ a symmetry of ${\mathbb{R}}^{3}$, as above. Let ${\cal R}_{0}$ be a cubiculated region and ${\mathbf{t}}_{0}$ be a slab tiling of ${\cal R}_{0}$. Let ${\cal R}_{1}=q[{\cal R}_{0}]$, a cubiculated region, and ${\mathbf{t}}_{1}=q[{\mathbf{t}}_{0}]$, a tiling of ${\cal R}_{1}$. Let $\kappa_{0}$ be a pair of colors and $\kappa_{1}=q[\kappa_{0}]$. Let ~$\tilde{\cal R}_{0,\kappa_{0}}$ and ~$\tilde{\cal R}_{1,\kappa_{1}}$ be the respective transformed regions, with domino tilings $\tilde{\mathbf{t}}_{0,\kappa_{0}}$ and $\tilde{\mathbf{t}}_{1,\kappa_{1}}$. We then have
$\operatorname{Tw}_{\kappa_{0}}({\mathbf{t}}_{0})=\operatorname{Tw}(\tilde{\mathbf{t}}_{0,\kappa_{0}})=\operatorname{sign}(q)\operatorname{Tw}(\tilde{\mathbf{t}}_{1,\kappa_{1}})=\operatorname{sign}(q)\operatorname{Tw}_{\kappa_{1}}({\mathbf{t}}_{1}).$
Proof.
The proof is easy and similar to that of Proposition 5.2. The reader may want to check consistency between the two propositions in a few cases. ∎
6 Connectivity under local moves
In this section we return to our main problem: connectivity under local moves of spaces of tilings. We first show that the space of slab tilings is connected via flips for a few special regions. We then show that there exists no finite fixed set of local moves which works for slab tilings of arbitrary boxes. Equivalently, there is no bound on the number of slabs which have to participate in each move.
We begin with two-floored regions.
Lemma 6.1.
Given a quadriculated disk ${\cal D}\subset\mathbb{R}^{2}$, let ${\cal R}={\cal D}\times[0,2]$, the three dimensional region with base ${\cal D}$ and two floors. The space of slab tilings of ${\cal R}$ is either empty or connected under flips.
Proof.
Let $\tilde{{\cal D}}\subset{\mathbb{R}}^{2}$ be a bounded region which can be tiled by $2\times 2$ squares (the region $\tilde{\cal D}$ need not be connected). We claim that there is only one way of tiling $\tilde{{\cal D}}$ with such squares. The proof of the claim is by induction on the number of unit squares of $\tilde{{\cal D}}$. The case where $\tilde{{\cal D}}$ comprises four unit squares is obvious. For the inductive step, notice there is only one way of placing a slab on a particular corner of $\tilde{{\cal D}}$, and then remove the four corresponding unit squares.
We return to ${\cal R}$. We may speak in terms of the floor diagram. On the first floor, omit all dominoes (i.e., parts of vertical slabs). The remaining planar region $\tilde{\cal D}_{1}$ must be tiled by $2\times 2$ squares, which can only be done in one way. On the second floor, omit all dominoes again. The remaining planar region ${\cal D}_{2}$ is the same as ${\cal D}_{1}$. Hence, it must be tiled the same way by $2\times 2$ squares.
A square on the first floor can be paired with the square which is right above it on the second floor. Flipping each one of these pairs creates a tiling with all slabs vertical, or, equivalently, identical domino-only configurations on both floors. By the domino flip connectedness result in dimension $2$ [5, 8], we can reach a given domino tiling on both floors, proving connectivity. ∎
Proposition 6.2.
Consider a box of the form $4\times 4\times N$, $N\in{\mathbb{N}}$. The space of slab tilings of the box has only one flip component.
Proof.
This can be done by casework and essentially consists of deducing the tree of possibilities of the tilings for each floor. This is similar to the computation of domino groups in [6]. We omit the computation, which is detailed in [9]. ∎
Figure 20: A tiling which admits no flips; $\operatorname{Tw}_{\kappa}=+4$ for $\kappa$ equal to red-green
Let us recall a fact about domino tilings and flips: there are often connected components of size 1, i.e., tilings which admit no flip. Figure 1 shows an example for the $3\times 3\times 2$ box; there are similar examples in larger boxes (see [6] for examples).
An example of a box whose space of slab tilings has components of size one (under flips) is the $6\times 6\times 5$ box. One such tiling is shown in Figure 20. Another component (for the same box) can be obtained by reflecting the tiling about the plane of the paper. This example admits slight variations for other regions. We can add a floor comprising only slabs before the first floor, or we can pile copies of the tiling of the above figure. Also, we can generalize the pattern for regions of the form $2N\times 2N\times 5$, as in Figure 21 for the case $8\times 8\times 5$.
Figure 21: Another tiling which admits no flips; in this case, $\operatorname{Tw}_{\kappa}=+12$
Recall that it is conjectured that the space of domino tilings of a 3d box is connected via flips and trits. We now show that there is no fixed finite set of local moves under which the set of slab tilings of arbitrary boxes becomes connected. We first study the family of examples in Figures 20 and 21.
Lemma 6.3.
Consider the region ${\cal R}=[0,2k]\times[0,2k]\times[0,5l]$, $k\geq 3$, $l\geq 1$ and a slab tiling ${\mathbf{t}}$ following the pattern in Figure 20 (piling copies if $l>1$). Now consider ${\cal R}_{0}=[0,2(k-1)]\times[0,2(k-1)]\times[0,5l]$. In ${\cal R}$, there is only one way of removing and placing back the slabs which are contained in the interior of ${\cal R}_{0}$.
Proof.
Refer to Figure 22. The region ${\cal R}_{0}\subset{\cal R}$ is bounded by the red curve. The region ${\cal R}_{1}=[0,2(k-2)]\times[0,2(k-2)]\times[0,5l]\subset{\cal R}_{0}$ is bounded by the dotted curve. We proceed by induction on $k$. We show the inductive process for five consecutive floors. The slab numbered $1$ is clearly stuck since it is “sandwiched” between two other slabs. Using this argument, on the first three floors, we proceed to deduce the positions of slabs which intersect the columns (in the floor diagram) between the dotted and red lines. In the figure, this corresponds to the slabs from $1$ to $8$. We repeat this, but now for slabs intersecting the rows between the dotted and red lines. This corresponds to slabs from $9$ to $14$.
1212$\scriptstyle$ 1212
88$\scriptstyle$ 88
66$\scriptstyle$ 66
77$\scriptstyle$ 77
1212$\scriptstyle$ 1212
1313$\scriptstyle$ 1313
1010$\scriptstyle$ 1010
22$\scriptstyle$ 22
44$\scriptstyle$ 44
66$\scriptstyle$ 66
1414$\scriptstyle$ 1414
1313$\scriptstyle$ 1313
11$\scriptstyle$ 11
33$\scriptstyle$ 33
55$\scriptstyle$ 55
1414$\scriptstyle$ 1414
1616$\scriptstyle$ 1616
1818$\scriptstyle$ 1818
2020$\scriptstyle$ 2020
$2$$\scriptstyle$ $2$
1515$\scriptstyle$ 1515
1717$\scriptstyle$ 1717
1919$\scriptstyle$ 1919
2121$\scriptstyle$ 2121
2020$\scriptstyle$ 2020
$1$$\scriptstyle$ $1$
99$\scriptstyle$ 99
Figure 22: The tiling from Figure 21, with labels
Now, for the last two floors, we first deduce the slabs intersecting the rows between the dotted and red lines ($15$ to $19$) and then we deduce the slabs intersecting the columns between the dotted and red lines ($20$ to $22$). We then apply the inductive step on the box bounded by the dotted curve. ∎
Corollary 6.4.
Let ${\cal R}$ and ${\mathbf{t}}$ be as in Lemma 6.3. Let $\tilde{\mathbf{t}}\neq{\mathbf{t}}$ be another slab tiling of ${\cal R}$. Then there are at least $32k-$ 3 slabs in ${\mathbf{t}}$ which are not in $\tilde{\mathbf{t}}$.
Proof.
Consider tilings ${\mathbf{t}}$ and $\tilde{\mathbf{t}}$ as in the statement. Let ${\cal X}\subseteq{\cal R}$ be the union of slabs in ${\mathbf{t}}$ which are not in $\tilde{\mathbf{t}}$. Alternatively, ${\cal X}$ is the union of slabs in $\tilde{\mathbf{t}}$ which are not in ${\mathbf{t}}$. Let $d$ be the number of slabs in ${\cal X}$. We need to prove that $d\geq 2k-3$.
We may assume without loss that ${\cal X}$ has connected interior. Indeed, otherwise let $\hat{\cal X}\subsetneqq{\cal X}$ be the closure of one of the connected components of the interior of ${\cal X}$. The region $\hat{\cal X}$ is tiled by ${\mathbf{t}}$ and $\tilde{\mathbf{t}}$ in two different ways. Let $\hat{\mathbf{t}}$ coincide with $\tilde{\mathbf{t}}$ in $\hat{\cal X}$ and with ${\mathbf{t}}$ elsewhere. The new tiling $\hat{\mathbf{t}}$ is a better example than $\tilde{\mathbf{t}}$ (i.e., with a smaller value of $d$).
As in Lemma 6.3 let ${\cal L}_{\operatorname{NE}}={\cal R}\smallsetminus{\cal R}_{0}\subset{\cal R}$ be the L shaped region to the NE direction. Define similar subsets ${\cal L}_{\operatorname{SE}}$, ${\cal L}_{\operatorname{SW}}$ and ${\cal L}_{\operatorname{NW}}$. By Lemma 6.3, we have ${\cal X}\cap{\cal L}_{\operatorname{NE}}\neq\varnothing$; the intersections of ${\cal X}$ with the other three L shaped sets are likewise nonempty. Let ${\cal M}_{\operatorname{North}}={\cal L}_{\operatorname{NE}}\cap{\cal L}_{\operatorname{NW}}$; define similarly ${\cal M}_{\operatorname{South}}$, ${\cal M}_{\operatorname{East}}$ and ${\cal M}_{\operatorname{West}}$. We claim that at least one of the two conditions below hold:
${\cal X}\cap{\cal M}_{\operatorname{North}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{South}},\qquad{\cal X}\cap{\cal M}_{\operatorname{East}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{West}}.$
Indeed, assume for instance that ${\cal X}\cap{\cal M}_{\operatorname{West}}=\varnothing$: the conditions ${\cal X}\cap{\cal L}_{NW}\neq\varnothing$ and ${\cal X}\cap{\cal L}_{SW}\neq\varnothing$ imply ${\cal X}\cap{\cal M}_{\operatorname{North}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{South}}$; the other cases are similar.
Assume without loss of generality that ${\cal X}\cap{\cal M}_{\operatorname{North}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{South}}$. Consider the planes $p_{x}=\{x\}\times[0,2k]\times[0,5l]$, $22\leq x\leq 2k-$ 2, which appear as horizontal lines in Figures 20 and 21. Given $x$, search for the slabs in ${\mathbf{t}}|_{{\cal X}}$ or $\tilde{\mathbf{t}}|_{{\cal X}}$ which cross $p_{x}$. We claim there are at least two such slabs. Notice first that if there are zero such slabs then ${\cal X}$ becomes essentially disconnected and we have a contradiction as above. Suppose then that there exist exactly one such slab. The region ${\cal X}_{\operatorname{North}}={\cal X}\cap([0,x]\times[0,2k]\times[0,5l])$ is tiled (by slabs, by either ${\mathbf{t}}$ or $\tilde{\mathbf{t}}$): the volume $V$ of ${\cal X}_{\operatorname{North}}$ is therefore a multiple of $4$. The region minus two unit cubes is also tiled (by either $\tilde{\mathbf{t}}$ or ${\mathbf{t}}$) and therefore $V-2$ is also a multiple of $4$, a contradiction. This completes the proof of the claim.
The total number $2d$ of slabs of both ${\mathbf{t}}$ and $\tilde{\mathbf{t}}$ in ${\cal X}$ is therefore at least $2(2k-3)$, completing the proof. ∎
Remark 6.5.
The corollary above is not sharp, the best estimate appears to be $12k-$ 1 (for all $l\geq 1$). The above formulation is sufficient for our main objective: no fixed set of local moves (involving a bounded number of slabs) connects all slab tilings of large boxes. More than that: given a fixed set of local moves, for sufficiently large $k$ the tiling ${\mathbf{t}}$ above admits none of these moves. $\diamond$
7 Possible values of the triple twist
Let ${\cal R}$ be a region. We denote $\operatorname{TTw}[{\cal R}]=\{\operatorname{TTw}({\mathbf{t}})|{\mathbf{t}}\text{ is a slab tiling of }{\cal R}\}$; the pairs $\kappa_{x},\kappa_{y},\kappa_{z}$ are implicit. Also, for a pair of colors $\kappa$, set $\operatorname{Tw}_{\kappa}[{\cal R}]=\{\operatorname{Tw}_{\kappa}({\mathbf{t}})|{\mathbf{t}}\text{ is a slab tiling of }{\cal R}\}$. The remainder of this section establishes bounds on the number of values the triple twist assumes in a given region, implying a bound on the number of slab flip connected components. The proof is inspired by a related construction in [4].
Lemma 7.1.
There exist constants $C_{+}>0$ and $C_{1}>0$ such that the following assertions hold: If ${\cal R}=[0,N]^{3}$ is a cube with even side lengths then $|\operatorname{TTw}({\mathbf{t}})|<C_{+}N^{4}$ for any slab tiling ${\mathbf{t}}$ of ${\cal R}$. Also for ${\cal R}=[0,N]^{3}$ we have $\lvert\operatorname{TTw}[{\cal R}]\rvert\leq C_{1}N^{12}$.
Proof.
Given a pair of colors $\kappa$, we bound $\lvert\operatorname{Tw}_{\kappa}[{\cal R}]\rvert$. The transformed region $\tilde{{\cal R}}$ is contained in a cube of side $N$. See Figures 18 and 23 for an example of the transformed region for $N=4$. In Figure 23, the floor is bounded by the dotted $4\times 4$ square. We show $3$ dominoes on the central column. In this example, $d_{y}=d_{z}=1$ and there is one effect.
$y$
$x$
Figure 23: A typical floor after transforming a $4\times 4\times N$ box
Fix $f$ a floor of $\tilde{{\cal R}}$, fix a column $c$ of $f$, and use axes as in Figure 23. Consider dominoes with a square in $c$: there are $d_{y}$ dominoes in the direction of the $y$-axis and $d_{z}$ dominoes in the direction of the $z$-axis. The number of possible effects happening within $c$ is at most $d_{y}d_{z}$. Since $d_{y}+d_{z}\leq N$, we have $d_{y}d_{z}\leq\frac{N^{2}}{4}$ by AM-GM inequality.
Now, varying $c$ and then $f$, we see there are at most $\frac{N^{4}}{4}$ effects within $\tilde{{\cal R}}$. Thus, for every tiling ${\mathbf{t}}$ we have $|\operatorname{Tw}_{\kappa}({\mathbf{t}})|\leq\frac{N^{4}}{16}$. The first claim therefore holds for $C_{+}=\frac{1}{4}$.
Since $N\geq 2$, we have
$\lvert\operatorname{Tw}_{\kappa}[{\cal R}]\rvert\leq\frac{N^{4}}{8}+1\leq\frac{5}{32}N^{4}.$
Since the above estimate holds for $\kappa_{x}$, $\kappa_{y}$ and $\kappa_{z}$, we obtain
$\lvert\operatorname{TTw}[{\cal R}]\rvert\leq\left(\frac{5}{32}\right)^{3}N^{12}\leq\frac{1}{2^{8}}N^{12}.$
This completes the proof, with $C_{1}=\frac{1}{2^{8}}$. ∎
Lemma 7.2.
Let ${\cal R}=[0,6n]\times[0,6n]\times[0,8n]$ be a box. For any $u,v\in[-n^{2},n^{2}]\cap{\mathbb{Z}}$ we construct a slab tiling ${\mathbf{t}}_{u,v}$ of ${\cal R}$ such that $\operatorname{TTw}({\mathbf{t}}_{u,v})=(0,0,2uv)$.
Proof.
Consider the subdivision of the $[0,6n]\times[0,6n]\times[0,8n]$ box in parts $\alpha,\beta$ and $\gamma$ as in Figure 24. We tile $\alpha$ with horizontal slabs. We can tile $\beta$ and $\gamma$ both as collections of $2n^{2}$ annuli. In Figure 24, there are four kinds of floors. Each of them appears $2n$ times. Each rectangle on each floor has dimensions either $6n\times 2n$, $2n\times 6n$ or $2n\times 2n$.
$y$
$x$
$\alpha$
$\alpha$
$\beta$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\beta$
$\beta$
$\gamma$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\gamma$
$\gamma$
$\beta$
$\alpha$
$\alpha$
$\gamma$
Figure 24: Floor diagram of subdivision of the $[0,6n]\times[0,6n]\times[0,8n]$ box
Consider the surfaces $S_{\beta}$ and $S_{\gamma}$ which are both squares of side $2n$, depicted as red cuts in Figure 25. Note that the top left vertex of a floor on the upper (resp. lower) diagram is of the form $(0,*,0)$ (resp. $(*,0,0)$). The surfaces $S_{\beta}$ and $S_{\gamma}$ are depicted in red.
$S_{\gamma}$
$z$
$x$
$S_{\beta}$
$y$
$z$
Figure 25: Example of $\beta$ ($y$-axis view) and $\gamma$ ($x$-axis view) on the $[0,12]\times[0,12]\times[0,16]$ box
Look at $S_{\beta}$ from the $y$-axis view. We regard each floor of $S_{\beta}$ with only grey pieces (a grey floor) as a domino tiling of an annulus, and we consider the flux with respect to the red cut. The flux across $S_{\beta}$ is the sum of the fluxes on all grey floors. The flux across $S_{\gamma}$ is defined analogously. We see the fluxes across each of the surfaces $S_{\beta}$ and $S_{\gamma}$ are integers in the interval $[-n^{2},n^{2}]$. We can choose the fluxes across $S_{\beta}$ and $S_{\gamma}$ to be equal to $u$ and $v$, respectively. We call this tiling ${\mathbf{t}}={\mathbf{t}}_{u,v}$.
We now verify ${\mathbf{t}}$ satisfies the desired properties. We will first see $\operatorname{Tw}_{\kappa_{x}}({\mathbf{t}})=\operatorname{Tw}_{\kappa_{y}}({\mathbf{t}})=0$. Without loss, we prove the first case. Consider ${\mathbf{t}}_{\kappa_{x}}$ and refer to Figure 14. After a sequence of flips, the slabs of $\alpha$ can be made horizontal. Slabs from $\beta$ and $\gamma$ from floors from $(2n+1)$-th to $(4n)$-th, that is, where $\beta$ passes through $\gamma$, are transformed into vertical dominoes in $\tilde{\mathbf{t}}_{\kappa_{x}}$. These vertical dominoes do not affect the ones from $\alpha$. Now, for $1\leq j\leq 2n$, the floors $j$ and $6n+1-j$ in $\tilde{\mathbf{t}}_{\kappa_{x}}$ are equal, except for the vertical dominoes, which have different colors on the floor diagram. This means the effects on floors $j$ and $6n+1-j$ cancel. Thus, $\operatorname{Tw}_{\kappa_{x}}({\mathbf{t}})=0$.
We now prove $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 2uv$. We will consider $\tilde{{\mathbf{t}}}$ as a dimer covering ${\mathbf{t}}^{*}$ on the graph ${\cal R}^{*}$ whose vertices are centers of squares in $\tilde{{\cal R}}$ and edges join centers of adjacent squares. We draw ${\cal R}^{*}$ on top of ${\cal R}$ for ease of visualization. We put red dots on the squares of ${\cal R}$ painted with good colors. We proceed to define surfaces $\tilde{S}_{\beta}$ and $\tilde{S}_{\gamma}$ as follows. From the $z$-axis view, look at the first floor of $\beta$ which intersects $S_{\beta}$. Consider the red dots immediately above or below the red cut and draw a polygonal line joining them from left to right as in Figure 26. Then, $\tilde{S}_{\beta}$ is a curtain with base this polygonal line and height $12n-$ 1. The definition of the surface $\tilde{S}_{\gamma}$ is analogous: we now consider red dots to the left or right of the red cut, also in Figure 26. In Figure 26, the polygonal lines of $\tilde{S}_{\beta}$ and $\tilde{S}_{\gamma}$ are dotted and the surfaces $S_{\beta}$ and $S_{\gamma}$ are represented as full lines.
$x$
$y$
Figure 26: Construction of the surfaces $\tilde{S}_{\beta}$ (above) and $\tilde{S}_{\gamma}$ (below) following the example in Figure 25
We see that if we twist one annulus of $\beta$ (resp $\gamma$) the flux across $S_{\beta}$ (resp $S_{\gamma}$) of the domino tiling in Figure 25 changes by ±$\pm 1$. Thus, the flux across $\tilde{S}_{\beta}$ (resp $\tilde{S}_{\gamma}$) changes by ±$\pm 2$. Note that translating the new surface cuts does not change the respective fluxes. We do not need to be careful with orientations and signs here. Also, $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})$ changes by ±$\pm 2v$ (resp ±$\pm 2u$).
To prove this, without loss, let ${\mathbf{t}}_{0}$ be ${\mathbf{t}}$ after twisting one of the annuli of $\beta$. The difference $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ has one non trivial cycle which bounds an oriented surface $\tilde{S}$. We draw $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ on top of ${\cal R}$ as in Figure 27. The surface $\tilde{S}$ bounded by $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ is drawn as black polygonal lines. Full lines represent the first and last floors of this cycle. The first and last floors of two annuli from $\tilde{\gamma}$ are shown as purple polygonal lines. Full segments represent the actual dimers in the tiling. Notice $\tilde{S}$ is tangent to $\tilde{S}_{\gamma}$. To calculate $t=\operatorname{Tw}(\tilde{\mathbf{t}}_{0})-\operatorname{Tw}(\tilde{{\mathbf{t}}})$, we must count dominoes from $\tilde{{\mathbf{t}}}$ passing through $S$. These are exactly the ones which are part of an annulus in $\tilde{\gamma}$. By twisting one annulus from $\tilde{\gamma}$ we see $t$ changes by ±$\pm 2$, hence $t=\pm 2v$ by induction.
Figure 27: Some floors of the $[0,12]\times[0,12]\times[0,16]$ box
$S_{\gamma}$
$S_{\beta}$
Figure 28: We continue with the example of Figure 25
Finally, we claim that if either flux across $\tilde{S_{\beta}}$ or $\tilde{S}_{\gamma}$ is equal to 00, we can take ${\mathbf{t}}$ to a tiling with only horizontal slabs, and thus with no twist. This means by twisting annuli of ${\mathbf{t}}$ to end in the above zero flux situation we end up proving $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 2uv$.
Without loss of generality we treat the first case of the above claim. Refer to Figure 28: on the first row we make the slabs in $\beta$ horizontal, on the second row we make the slabs in $\gamma$ and its hole horizontal. If this holds, then we can flip the slabs in $\beta$ so that all of them are horizontal. In particular, the slabs passing through $\gamma$’s hole are horizontal. Look at $\gamma$ from the $x$-axis view. For every pair of consecutive floors of $\gamma$ we can make all slabs horizontal by using the domino flip connectedness result for planar regions [5, 8]. ∎
Lemma 7.3.
Let ${\cal R}=[0,16n]^{3}$ be a box. For any triple $t\in([-2n^{4},2n^{4}]\cap 2{\mathbb{Z}})^{3}$ we construct a slab tiling ${\mathbf{t}}$ of ${\cal R}$ such that $\operatorname{TTw}({\mathbf{t}})=t$.
Proof.
For numbers $u_{i,j},v_{i,j}\in[-n^{2},n^{2}]$, $1\leq i\leq 2$, $1\leq j\leq 3$ we can construct tilings ${\mathbf{t}}_{i,j}$ of the $[0,6n]\times[0,6n]\times[0,8n]$ box such that $\operatorname{TTw}({\mathbf{t}}_{j})=2u_{i,j}v_{i,j}\mathbf{e_{j}}$ by Lemma 7.2 (here $\mathbf{e_{j}}\in{\mathbb{Z}}^{3}$ is the $j$-th vector in the canonical base). After doing this, we divide the $[0,16n]^{3}$ box in $8$ cubes of side $8n$. Each of these smaller cubes is tiled using one of the ${\mathbf{t}}_{i,j}$ on one corner, and horizontal slabs to fill the rest of the space. Two of the smaller cubes will receive only horizontal slabs. This new tiling is named ${\mathbf{t}}$.
Let $t=(t_{1},t_{2},t_{3})$. Choose
$u_{1,j}=n^{2},\quad v_{1,j}=\left\lfloor\frac{t_{j}}{2n^{2}}\right\rfloor,\quad u_{2,j}=1,\quad v_{2,j}=n^{2}\left\{\frac{t_{j}}{2n^{2}}\right\}.$
Here, for $s\in{\mathbb{R}}$, we have $s=\lfloor s\rfloor+\{s\}$ where $\lfloor s\rfloor\in{\mathbb{Z}}$ and $\{s\}\in[0,1)$. We have $t_{j}=2u_{1,j}v_{1,j}+2u_{2,j}v_{2,j}$, as desired. Then, for each $j$, $\operatorname{TTw}({\mathbf{t}}_{1,j}+{\mathbf{t}}_{2,j})=t_{j}\mathbf{e_{j}}\in([-2n^{4},2n^{4}]\cap 2{\mathbb{Z}})^{3}$, where ${\mathbf{t}}_{1,j}+{\mathbf{t}}_{2,j}$ is the tiling of the two cubes containing ${\mathbf{t}}_{1,j}$ and ${\mathbf{t}}_{2,j}$. Summing the above for all $1\leq j\leq 3$ yields $\operatorname{TTw}({\mathbf{t}})=(t_{1},t_{2},t_{3})$. ∎
Lemma 7.4.
There is a constant $C_{0}>0$ such that the following assertion holds: If ${\cal R}=[0,N]^{3}$, $N\geq 256$, is a cube with even side lengths (and thus slab tileable) then $\lvert\operatorname{TTw}[{\cal R}]\rvert\geq C_{0}N^{12}$.
Proof.
Write $N=16M+2k$ for some $0\leq k\leq 7$. We can divide ${\cal R}$ into two parts ${\cal R}_{1}$ and ${\cal R}_{2}$ as in Figure 29 (there is one kind of floor: it comprises one $16M\times 16M$ square on the lower left corner, ${\cal R}_{1}$, and its complement, ${\cal R}_{2}$). For a given $\mathbf{v}\in([-2M^{4},2M^{4}]\cap 2{\mathbb{Z}})^{3}$, we tile ${\cal R}_{1}$ with a tiling ${\mathbf{t}}_{1}$ having $\operatorname{TTw}({\mathbf{t}}_{1})=\mathbf{v}$ and we tile ${\cal R}_{2}$ with horizontal slabs.
$\mathcal{R}_{1}$
$\mathcal{R}_{2}$
Figure 29: Diagram of subdivision of ${\cal R}$
Then, $\lvert\operatorname{TTw}[{\cal R}]\rvert\geq(2M^{4})^{3}=8M^{12}=8(\frac{N-2k}{16})^{12}=c(N-2k)^{12}$, where $c=\frac{1}{2^{45}}$. We seek $C>0$ such that $(N-2k)^{12}\geq CN^{12}$ holds. This happens if and only if $C\leq(1-\frac{2k}{N})^{12}$. By minimizing the RHS with the constraints $N\geq 256$ and $k\leq 7$, we can take $C=\frac{1}{2}\leq(1-\frac{14}{256})^{12}$. Finally, we ca ${\cal R}=[0,N]^{3}$n take $C_{0}=cC=\frac{1}{2^{46}}$. ∎
We summarize the previous lemmas in the proposition below.
Proposition 7.5.
For $C_{1}=\frac{1}{2^{8}}$ and $C_{0}=\frac{1}{2^{46}}$, the following assertion holds:
If $N\geq 256$ is an even integer, then
$C_{0}N^{12}\leq\lvert\operatorname{TTw}[[0,N]^{3}]\rvert\leq C_{1}N^{12}.$
Proof.
We take $C_{0}$ and $C_{1}$ as in the proofs of Lemmas 7.1 and 7.4. ∎
We are also ready for Theorem 1.
Proof of Theorem 1.
We use the max norm and give explicit values for $C_{\pm}$ assuming $N\geq 256$. As in Lemma 7.1, we can take $C_{+}=\frac{1}{4}$. We claim that $C_{-}=1/{2^{16}}$ works. Indeed, take $n=\left\lfloor\frac{N}{16}\right\rfloor$ and consider the smaller cube ${\cal R}_{-}=[0,16n]^{3}$. Notice that $N^{4}<2^{17}n^{4}$. Each coordinate of $t$ is an even integer with absolute value smaller than $C_{-}N^{4}<2n^{4}$. Apply Lemma 7.3 to deduce that there exists a tiling ${\mathbf{t}}_{-}$ of ${\cal R}_{-}$ with $\operatorname{TTw}({\mathbf{t}}_{-})=t$. Finally, construct a tiling ${\mathbf{t}}$ of ${\cal R}=[0,N]^{3}\supseteq[0,16n]^{3}$ by starting with ${\mathbf{t}}_{-}$ and completing the remaining region with horizontal slabs. We have $\operatorname{TTw}({\mathbf{t}})=\operatorname{TTw}({\mathbf{t}}_{-})=t$, as desired. ∎
8 Final remarks
There are many aspects of the theory of slab tilings that we seek to understand better. We list a few.
The first is that in proposition 7.5, the bounding constants were not optimised and thus are far from being sharp.
The second is that we would like to study the flip connected components in more classes of examples, not just the ones in proposition 6.2. For example, we do not know what happens in boxes of height $3$, but we guess there is only one component.
The third is regarding probabilistic results such as the ones for domino tilings in [7]. For example, for slab tilings we would like to know the distribution for the sizes of connected components of large regions.
The fourth is regarding other moves besides the flip. In the domino case, there is also the trit, and the question on whether all tilings of a given region are connected by flips and trits is open. In the slab case, the existence of a set of local moves which connect all or most tilings of a given region is hindered by Remark 6.5. In the example of Figure 21, however, it is possible to perform moves involving a quantity of pieces proportional to the lengths of the sides of the box. It can also be asked how common is a situation comparable to the example in Lemma 6.3 and Corollary 6.4. It would be interesting to clarify the situation in greater generality.
Also, one property observed in the examples of slab tilings we studied so far concerns the parity of the entries of the triple twist. Our final and main conjecture is the following.
Conjecture 8.1.
Let ${\cal R}$ be a region tileable by slabs. Then there exist $t_{x},t_{y},t_{z}\in\{0,1\}\subset{\mathbb{Z}}$ with the following property: For every slab tiling ${\mathbf{t}}$ of ${\cal R}$, the integers
$\operatorname{Tw}_{\kappa_{x}}({\mathbf{t}})-t_{x},\quad\operatorname{Tw}_{\kappa_{y}}({\mathbf{t}})-t_{y},\quad\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})-t_{z}$
are all even. If ${\cal R}$ is a box then $t_{x}=t_{y}=t_{z}=0$.
Figure 30: A tiling ${\mathbf{t}}$ with $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 1$
Figure 30 shows a region for which the conjecture holds with $t_{z}=1$.
9 Declarations
9.1 Funding and Acknowledgements
The second author is thankful for the generous support of CNPq, CAPES and FAPERJ (Brazil). The first and third authors are thankful to Projeto Arquimedes and to CNPq for the support in the PIBIC undergraduate project.